OpenStudy (anonymous):

Identify the period, phase shift and vertical shift of y=3tan(x-pi/2)+1. (if applicable) I really just need the answer. Just being honest

OpenStudy (anonymous):

your equation is set up as y= A sin (Bx - c) + D a = Amplitude 2π/b = period c/b = phase shift d = vertical shift Sooooooo Amplitude = 3, period is 2π, PS = (π/2)/2 (simplify that, I'm lazy.) and vs = 1

OpenStudy (anonymous):

Oh my god you rock, wanna help me with some more? i'll ask on a different one so you can get another medal?

OpenStudy (anonymous):

I JUST learned that today XDDDDD

OpenStudy (anonymous):

My Alg2/Trig class ends Friday. I really don't know that much XD

OpenStudy (anonymous):

my AFM class ends Friday too!

OpenStudy (anonymous):

would you help me with some more?